# Dialogue Systems for Emotional Support via Value Reinforcement

### v2

@@ Abstract
Emotional support dialogue systems aim to reduce help-seekers’ distress and help them overcome challenges. While human values—core beliefs that shape an individual’s priorities—are increasingly emphasized in contemporary psychological therapy for their role in fostering internal transformation and long-term emotional well-being, their integration into emotional support systems remains underexplored. To bridge this gap, we present a value-driven method for training emotional support dialogue systems designed to reinforce positive values in seekers. Notably, our model identifies which values to reinforce at each turn and how to do so, by leveraging online support conversations from Reddit. We evaluate the method across support skills, seekers’ emotional intensity, and value reinforcement. Our method consistently outperforms various baselines, effectively exploring and eliciting values from seekers. Additionally, leveraging crowd knowledge from Reddit significantly enhances its effectiveness. Therapists highlighted its ability to validate seekers’ challenges and emphasize positive aspects of their situations—both crucial elements of value reinforcement. Our work, being the first to integrate value reinforcement into emotional support systems, demonstrates its promise and establishes a foundation for future research. [^1] All source code and data will be made publicly available upon its publication.
@@ 1 Introduction
Emotional support aims to help individuals (seekers) in addressing everyday emotional difficulties, such as relationship conflicts and workplace stress, by offering reassurance, acceptance, and encouragement (Atoum and Al-Shoboul, 2018; Burleson, 2003). Recent advancements in large language models have accelerated the development of dialogue systems designed to provide emotional support (supporters) (Deng et al., 2024; Zhang et al., 2023; Chen et al., 2023). Many models have focused on reinforcing positive emotions in seekers. However, emotional changes alone may not adequately capture deeper intrinsic transformations within the seeker, potentially reducing the long-term impact of emotional support (Blackledge and Hayes, 2001). For instance, a seeker’s perfunctory “Thank you”, used as a conversational pleasantry, receives a higher positivity score (0.758) from a sentiment classifier than the response shown in Figure 1 (0.583) [^2] We employed EmoLlama-Chat-7B (Liu et al., 2024), which demonstrates superior performance in this task.. The low positivity score of the latter response may be attributed to phrases like “I know it’s not going to be easy”, which could be perceived as negative. However, the latter response demonstrates a stronger commitment and willingness to change, highlighting the importance of evaluating support effectiveness beyond emotion alone.
@@
We conducted a comprehensive evaluation in terms of supporter capabilities, the seeker’s ultimate relief, and value reinforcement. The results demonstrate that our model outperforms most baselines in supporter capabilities and value reinforcement, while maintaining a competitive level of seeker relief. Notably, the model’s strength in value reinforcement is highlighted in evaluations by expert therapists [^3] All therapists mentioned in this paper refer to two licensed clinical psychologists with over three years of clinical experience, who are also co-authors of this paper.. Specifically, it excels at effectively validating the seeker’s challenges and emphasizing positive aspects of the seeker’s situation, which form the foundation of value reinforcement. These results highlight that value reinforcement is a promising direction for future research.
@@ Results.
These findings demonstrate that value reinforcement in seekers positively impacts the effectiveness of emotional support and motivate our research approach to designing dialogue systems that aim to reinforce seekers’ values.
@@ 5 Method
The overall framework, illustrated in Figure 3, consists of three core components: (1) target value detector identifies values to reinforce at each turn; (2) reference generator produces utterances to effectively promote these values from the seeker; (3) supporter model determines strategies and generates responses based on the identified target values and the reference responses. To ensure the reproducibility of our work, we will release our source code, trained models, and simulation data publicly.
@@ 5.2 Reference Generator
This stage aims to enhance the SFT model’s generation quality through DPO. The preference dataset is constructed as follows. Given a dialogue history ($o_{1},c_{1},o_{2},c_{2},...,c_{t-1},o_{t}$), the original supporter comment $c_{t}$ is designated as the chosen response, as it successfully promoted the target values $v_{t+1}$. The rejected response is selected as another comment to $o_{t}$, denoted by $c_{t}^{\prime}$, randomly sampled from the siblings of $c_{t}$ (i.e., other comments under the same dialogue history). $c_{t}^{\prime}$ is a natural response to the dialogue history but is likely suboptimal for promoting the target values $v_{t+1}$ originally promoted by $c_{t}$. To mitigate the risk that $c_{t}^{\prime}$ is inadvertently effective for promoting $v_{t+1}$, we exclude any overlapping values between $v_{t+1}$ and $v_{t+1}^{\prime}$ (i.e., the seeker’s values expressed in $o_{t+1}^{\prime}$ in response to $c_{t}^{\prime}$), retaining up to three distinct target values unique to the chosen response in the final preference dataset. Detailed training methods and results are provided in Appendix D.2.
@@ 5.3 Supporter Model
SFT requires large-scale emotional support conversations grounded in value reinforcement. To obtain such data, we opt to use dialogue simulation with a seeker simulator (Section 5.4). We use GPT-4o-mini for both the supporter and seeker simulators to generate data for training a smaller model [^5] In our pilot experiment, zero-shot Llama-3-8B-Instruct was found to be unsuitable as a supporter simulator due to issues like repetitive responses and biases in reference usage.. The simulators engage in interactions by iteratively producing an utterance based on the ongoing dialogue history as a prompt and appending it to the history prompt.
@@
We construct the preference data as follows. For each dialogue, every supporter turn is assumed to have two response candidates (i.e., one with and one without using the reference response). We compute the expected reward for each response to determine the chosen and rejected responses for DPO. This reward is based on how well the intended target values at turn $t$ are reflected in the seeker’s subsequent utterances. The reward for a supporter response at turn $t$, $u_{t}^{\text{sup}}$, is calculated as:
$R(u_{t}^{\text{sup}})=\sum_{k=1}^{h}\gamma^{k-1}N_{t+k}(v_{t+1})$ (3)
where $N_{t+k}(v_{t+1})$ is the frequency of the values targeted at turn $t$ appearing in the seeker’s utterance at turn $t+k$, $h$ is the look-ahead horizon (the number of future steps considered), and $\gamma$ is a discount factor balancing immediate and future rewards. A response pair is added to the DPO dataset only if their reward difference exceeds a threshold $T_{\text{diff}}$.
@@ 5.4 Seeker Simulator
The seeker simulator is based on GPT-4o-mini, with its detailed design and validity provided in Appendix  E. To summarize, we extensively verified the quality of the seeker simulator using human and automated evaluations. Human evaluators judged our seeker simulator to be as natural as, or more natural than, human seekers (Figure 4). Additionally, it produced utterances that more closely resemble human seekers’ responses in content, emotional tone, and value alignment, compared to other simulators.
@@ 6.1 Evaluation Methods
We evaluate various supporter models through conversations with the seeker simulator, using the 120 held-out seeker personas for testing. A conversation is considered complete if the seeker simulator generates “[END]” or if the seeker’s emotion score, as calculated by EmoLlama-Chat-7B (Liu et al., 2024), reaches 0.6 or higher with gratitude expressions (e.g., “thank you”). Interactions are limited to a maximum of 20 turns, based on the average conversation length of 15 turns observed in the ESConv dataset. Only conversations concluding within this limit are included in the evaluation.
@@ 6.4.1 Effectiveness of Value Targeting and Reference Responses
Table 3: Comparison of models based on ES-Skills and ES-Intensity. ES-Value: The win-ratio of each model against Ours (DPO). Statistically significant differences compared to our DPO model are marked with * ($p$-value $<0.05$), and differences with $p$-value $<0.1$ are marked with $\dagger$, as determined by the Mann-Whitney U test.
@@ 6.4.3 Success of Target Value Reinforcement
As our approach focuses on setting target values and reinforcing them, we evaluated its effectiveness in reinforcing the target values. To this end, we compared our models with other prompt-based models that also take target values and reference responses as inputs.
@@ 6.4.4 Generalization beyond Reddit Data
Similarly, for ES-Skills (Table 3), DPO (Cactus) performed comparably to or better than the baselines, particularly excelling in emotional support metrics. However, it scored significantly lower on the Experience metric compared to the original DPO model. Since Reddit contains numerous shared personal narratives, the reference response generator appears to benefit more from Reddit than from Cactus, supporting our decision to use Reddit.
@@ 6.4.5 Expert Evaluation
To gain deeper insights into the value reinforcement capabilities of our supporter models, two licensed clinical psychologists with over three years of clinical experience conducted a qualitative analysis of dialogues generated by our DPO model.

### v3

@@ Abstract
Emotional support dialogue systems aim to reduce help-seekers’ distress and help them overcome challenges. While human values—core beliefs that shape an individual’s priorities—are increasingly emphasized in contemporary psychological therapy for their role in fostering internal transformation and long-term emotional well-being, their integration into emotional support systems remains underexplored. To bridge this gap, we present a value-driven method for training emotional support dialogue systems designed to reinforce positive values in seekers. Notably, our model identifies which values to reinforce at each turn and how to do so, by leveraging online support conversations from Reddit. We evaluate the method across support skills, seekers’ emotional intensity, and value reinforcement. Our method consistently outperforms various baselines, effectively exploring and eliciting values from seekers. Additionally, leveraging crowd knowledge from Reddit significantly enhances its effectiveness. Therapists highlighted its ability to validate seekers’ challenges and emphasize positive aspects of their situations—both crucial elements of value reinforcement. Our work, being the first to integrate value reinforcement into emotional support systems, demonstrates its promise and establishes a foundation for future research. [^1] Our code and dataset are available at https://github.com/holi-lab/ES-Value.
@@ 1 Introduction
Emotional support aims to help individuals (seekers) in addressing everyday emotional difficulties, such as relationship conflicts and workplace stress, by offering reassurance, acceptance, and encouragement (Atoum and Al-Shoboul, 2018; Burleson, 2003). Recent advancements in large language models have accelerated the development of dialogue systems designed to provide emotional support (supporters) (Deng et al., 2024; Zhang et al., 2023; Chen et al., 2023). Many models have focused on reinforcing positive emotions in seekers. However, emotional changes alone may not adequately capture deeper intrinsic transformations within the seeker, potentially reducing the long-term impact of emotional support (Blackledge and Hayes, 2001). For instance, a seeker’s perfunctory “Thank you”, used as a conversational pleasantry, receives a higher positivity score (0.758) from a sentiment classifier than the response shown in Figure 1 (0.583). [^2] We employed EmoLlama-Chat-7B (Liu et al., 2024), which demonstrates superior performance in this task. The low positivity score of the latter response may be attributed to phrases like “I know it’s not going to be easy”, which could be perceived as negative. However, the latter response demonstrates a stronger commitment and willingness to change, highlighting the importance of evaluating support effectiveness beyond emotion alone.
@@
We conducted a comprehensive evaluation in terms of supporter capabilities, the seeker’s ultimate relief, and value reinforcement. The results demonstrate that our model outperforms most baselines in supporter capabilities and value reinforcement, while maintaining a competitive level of seeker relief. Notably, the model’s strength in value reinforcement is highlighted in evaluations by expert therapists. [^3] All therapists mentioned in this paper refer to two licensed clinical psychologists with over three years of clinical experience, who are also co-authors of this paper. Specifically, it excels at effectively validating the seeker’s challenges and emphasizing positive aspects of the seeker’s situation, which form the foundation of value reinforcement. These results highlight that value reinforcement is a promising direction for future research.
@@ Results.
These findings support that value reinforcement in seekers positively impacts the effectiveness of emotional support and motivate our research approach to designing dialogue systems that aim to reinforce seekers’ values.
@@ 5 Method
The overall framework, illustrated in Figure 3, consists of three core components: (1) target value detector identifies values to reinforce at each turn; (2) reference generator produces utterances to effectively promote these values from the seeker; (3) supporter model determines strategies and generates responses based on the identified target values and the reference responses.
@@ 5.2 Reference Generator
This stage aims to enhance the SFT model’s generation quality through DPO. The preference dataset is constructed as follows. Given a dialogue history ($o_{1},c_{1},o_{2},c_{2},...,c_{t-1},o_{t}$), the original supporter comment $c_{t}$ is designated as the preferred response, as it successfully promoted the target values $v_{t+1}$. The rejected response is selected as another comment to $o_{t}$, denoted by $c_{t}^{\prime}$, randomly sampled from the siblings of $c_{t}$ (i.e., other comments under the same dialogue history). $c_{t}^{\prime}$ is a natural response to the dialogue history but is likely suboptimal for promoting the target values $v_{t+1}$ originally promoted by $c_{t}$. To mitigate the risk that $c_{t}^{\prime}$ is inadvertently effective for promoting $v_{t+1}$, we exclude any overlapping values between $v_{t+1}$ and $v_{t+1}^{\prime}$ (i.e., the seeker’s values expressed in $o_{t+1}^{\prime}$ in response to $c_{t}^{\prime}$), retaining up to three distinct target values unique to the preferred response in the final preference dataset. Detailed training methods and results are provided in Appendix D.2.
@@ 5.3 Supporter Model
SFT requires large-scale emotional support conversations grounded in value reinforcement. To obtain such data, we opt to use dialogue simulation with a seeker simulator (Section 5.4). We use GPT-4o-mini for both the supporter and seeker simulators to generate data for training a smaller model. [^5] In our pilot experiment, zero-shot Llama-3-8B-Instruct was found to be unsuitable as a supporter simulator due to issues like repetitive responses and biases in reference usage. The simulators engage in interactions by iteratively producing an utterance based on the ongoing dialogue history as a prompt and appending it to the history prompt.
@@
We construct the preference data as follows. For each dialogue, every supporter turn is assumed to have two response candidates (i.e., one with and one without using the reference response). We compute the expected reward for each response to determine the preferred and rejected responses for DPO. This reward is based on how many intended target values at that turn are expressed in the seeker’s subsequent utterances. The reward for a supporter response at turn $t$, $u_{t}^{\text{sup}}$, is:
$R(u_{t}^{\text{sup}})=\sum_{k=1}^{h}\gamma^{k-1}N_{t+k}$ (3)
where $N_{t+k}$ is the frequency of the values targeted at turn $t$ appearing in the seeker’s utterance at turn $t+k$, $h$ is the look-ahead horizon (the number of future steps considered), and $\gamma$ is a discount factor balancing immediate and future rewards. A response pair is added to the DPO dataset only if their reward difference exceeds a threshold $T_{\text{diff}}$.
@@ 5.4 Seeker Simulator
The seeker simulator is based on GPT-4o-mini, with its detailed design and validity provided in Appendix  E. To summarize, we extensively verified the quality of the seeker simulator using human and automated evaluations. Human evaluators judged our seeker simulator to be as natural as, or more natural than, human seekers (Figure 4). Additionally, utterances produced by GPT-4o-mini as the seeker simulator more closely resemble human seekers’ utterances in content, emotional tone, and value alignment, compared to other models.
@@ 6 Experiments
We evaluate our model ES-VR (Emotional Support via Value Reinforcement) through comprehensive experiments.
@@ 6.1 Evaluation Methods
We evaluate various supporter models through conversations with the seeker simulator using the 120 held-out seeker personas for testing. A conversation is considered complete if the seeker simulator generates “[END]” or if the seeker’s emotion score, as calculated by EmoLlama-Chat-7B (Liu et al., 2024), reaches 0.6 or higher with gratitude expressions (e.g., “thank you”). Interactions are limited to a maximum of 20 turns, based on the average conversation length of 15 turns observed in the ESConv dataset. Only conversations concluding within this limit are included in the evaluation.
@@ 6.4.1 Effectiveness of Value Targeting and Reference Responses
Table 3: Comparison of models based on ES-Skills and ES-Intensity. ES-Value: The win-ratio of each model against ES-VR (DPO). Statistically significant differences compared to our DPO model are marked with * ($p$-value $<0.05$), and differences with $p$-value $<0.1$ are marked with $\dagger$, as determined by the Mann-Whitney U test.
@@ 6.4.3 Success of Target Value Reinforcement
As our approach focuses on setting target values and reinforcing them, we evaluated its effectiveness in reinforcing the target values. To this end, we compared our ES-VR (DPO) model with other prompt-based models that also take target values and reference responses as inputs.
@@ 6.4.4 Generalization beyond Reddit Data
Similarly, for ES-Skills (Table 3), DPO (Cactus) performed comparably to or better than the baselines, particularly excelling in emotional support metrics. However, it scored significantly lower on the Experience metric compared to the original DPO model based on Reddit. Since Reddit contains numerous shared personal narratives, the reference response generator appears to benefit more from Reddit than from Cactus, supporting our decision to use Reddit.
@@ 6.4.5 Expert Evaluation
To gain deeper insights into the value reinforcement capabilities of our supporter models, two licensed clinical psychologists with over three years of clinical experience conducted a qualitative analysis of dialogues generated by our ES-VR (DPO) model.
@@ 7 Conclusion
Acknowledgments
This work was supported by the New Faculty Startup Fund and the Creative-Pioneering Researchers Program from Seoul National University. It was also supported by the National Research Foundation of Korea (NRF) grants funded by the Korean government (MSIT) (RS-2024-00333484, RS-2024-00455484).